Grade Upto college level Engineering Entrance Exams

An open organ pipe has a fundamental frequency of 300 Hz. The first overtone of this organ pipe is the same as the first overtone of a closed organ pipe. The length of the closed organ pipe is (a) 10 cm (b) 41 cm (c) 82 cm (d) 164 cm

sumit kumar
12 Years agoGrade Upto college level

2 Answers

Arun
6 Years ago
Given An open organ pipe has a fundamental frequency of 300 Hz. The first overtone of a closed organ pipe has the same frequency as the first overtone of the open pipe. Velocity of sound in air =330m/s. The length of the pipes are
Frequency is v/2L = 300 hz
So first overtone will be v/l = 600 hz
So frequency of 1st overtone of open pipe is equal to frequency of 1st overtone of closed pipe.
Speed of sound is 330 m/s
 Now 3v/4l = 600
 600 x 4l = 3v
 4l = 3v/600
 4l = 3v/600
l = 3 x 330/600 x 4
l = 33/80
l = 0.4125 m
l = 41.25 cm
length of pipe is 41.25 cm
 
Vikas TU
6 Years ago
Frequency is v/2L = 300 hz
So first overtone will be v/l = 600 hz
So frequency of 1st overtone of open pipe is equal to frequency of 1st overtone of closed pipe.
Speed of sound is 330 m/s
 Now 3v/4l = 600
 600 x 4l = 3v
 4l = 3v/600
 4l = 3v/600
l = 3 x 330/600 x 4
l = 33/80
l = 0.4125 m
l = 41.25 cm
 

question mark
Hello,
 
Can we learn and prepare for MHT-CET online ?
engineering entrance exams

Last Activity: 4 Years ago